A plano-convex lens (f = 20 cm) is silvered at plane surface. Now focal length will be :
(A) 20 cm (B) 40 cm (C) 30 cm (D) 10 cm

Grade:

2 Answers

Arun
3 years ago
 (1/f) = (n – 1) [(1 / R1) – (1 / R2)] , for plano convex lens , R1 = R, R2 = ∞ hence (1 / 20) = (1.5 – 1) [(1 / R) – (1 / ∞)] = (0.5 / R) ‑‑‑‑‑‑‑‑ (n = 1.5) ∴  R = 10 cm Also for rarer to denser medium , (n2 / v) – (n1 / u) = [(n2 – n1) / R] Here   n1 = 1,  n2 = 1.5,      v = f,  u = ∞ hence (1.5 / f) – (1 / ∞) = [(1.5 – 1) / 10]  ∴ (1.5 / f) = (0.5 / 10)  ∴ f = + 30 cm.
Vikas TU
3 years ago
A Plano convex lens is silvered at plane surface. It's focal length will be 10 cm. Hence, A Plano convex lens is silvered at plane surface. It's focal length will be 10 cm.